# Competition for binding targets results in paradoxical effects for simultaneous activator and repressor action - Extended Version

### v1

@@ I INTRODUCTION AND BACKGROUND
The field of synthetic biology has as its ultimate goal to program new or modify existing biological systems, for applications ranging from cell therapies and regenerative medicine to biosensing and biofuel production. In general, a significant obstacle to the development of synthetic biological circuits is the influence of compositional context: the fundamental characteristics of a circuit alter in the presence of additional components due to competition for resources, off-target interactions, genetic context, growth rate feedback loops, and retroactivity effects. See e.g. [1, 2, 3, 4, 5, 6] and [7] for an overview, Unless one designs mathematically-validated control circuits to compensate for uncertainty, designers will need to re-adjust each part whenever new elements are integrated into a system.
In general, a significant obstacle to the development of synthetic biological circuits is the influence of compositional context: the fundamental characteristics of a circuit alter in the presence of additional components due to competition for resources and retroactivity effects. Unless one designs mathematically-validated control circuits to compensate for uncertainty, designers need to re-adjust each part whenever new elements are integrated into a system.
@@
Since it is hard to test this mechanistic hypothesis through epigenetic modifications in mammalian cells, we propose here a synthetic biology analog involving transcription factors in bacterial cells. In this paper, we review the mechanism and make mathematical predictions from a model, as well as a proposed implementation using CRISPR/a technology. Experimental work is ongoing. In the remainder of this section, we review our previous results [8].
@@ I-A Review of previous work
Motivation: Paradoxical gene activity pattern. In the aforementioned experiments, a paradoxical pattern of gene activation was observed that cannot be easily explained by known local gene regulatory interactions. Consider a gene of interest (e.g, ZEB1 which is a major driver of EMT). As shown in Figure 2,
@@
the gene is nominally repressed due to the dominant activity of the repressor (PRC2). However, when the repressor is knocked out, the gene achieves a mediocre amount of activation. In the second knockout experiment, an activating regulator (KMT2D) is knocked out. In that case, the gene achieves maximal activation despite the fact that its main repressor is not knocked-out. The main question is: how would the direct knockout of a repressing regulator be less effective in activating a gene compared to knocking out an activator?. We summarize our answer next.
@@
    There is a large number of targets per regulator, e.g hundreds or thousands.
@@
This simplified model can partially explain the paradoxical results shown in Figure 2. It shows how repressor knockout can fail to activate a target gene, and how can an activator knockout activate the target. However, it does not show how can an activator knockout be more effective at activation than a repressor knockout. This is since the mechanism of activation depends on the sequestration of the repressor, hence it cannot yield an activation that is stronger than a full repressor knockout. We tackle this issue next.
The overall model. In order to recapture the full behavior in Figure 2, we need to add a second activator to the two factor model in Figure 4, as shown in Figure 5-a).
When the activator $A$ is knocked out and the repressor $R$ is sequestered by off-targets, the activator $A^{\prime}$ faces no longer any competition and can activate the target gene fully. A natural question might arise: when the repressor is knocked out in Figure 5-b), why cannot $A^{\prime}$ activate the target? The answer lies in the asymmetry in competition for off-targets. This model requires both $A,A^{\prime}$ to be sequestered when $R$ is knocked-out, but only $R$ being sequestered when $A$ is knocked out. This can be realized when each of them has different affinities for the target gene and the off-targets. Therefore, the overall model is depicted in Figure 5.
More generally, our model can admit all possible expressions levels for the three experimental scenarios. This is achieved by varying two parameters per regulator: a “local” association ratio, and a “global” association ratio, where the first describes the binding affinity of the regulator to a target of interest, while the second describes it binding affinity to off-targets. More details are available in [8].
Figure 5: The three-factor model that can explain paradoxical behaviors in gene expression upon knockouts.
As explained in the introduction, it is very difficult to test this mechanistic explanation of the paradoxical effects through epigenetic modifications in mammalian cells. Thus, in this work, we describe and analyze a synthetic biology analog that uses transcription factors in bacterial cells. We start with building the two-factor model, while the three factor model is planned for subsequent work.
@@ II Proposed synthetic construct
The proposed circuit (shown in Figure 6) consists of a pool of shared resources, activators, and repressors, regulating the production of the output protein by the target gene while competing with the rest of the genome, modeled as decoy sites.
@@
The transcription factors, activator ($A$, produced at a constant rate $u_{A}$) and repressor ($R$, produced at a constant rate $u_{R}$) bind to the target sites to transform into an active form ($T_{A}$) and silenced form ($T_{R}$). The active form of the target undergoes transcription and translation to produce the output protein at a rate $\kappa>\beta$, where $\beta$ is the basal transcription rate in the inactive form of the target. The decoy sites ($D$) sequester the resources by forming $D_{A}$ and $D_{R}$ complexes respectively. The chemical reactions involved are:
@@
The corresponding reaction rate equations (RREs) can be obtained using mass action kinetics as [11]:
@@
where $\gamma$ and $\bar{\delta}$ are the corresponding decay rates constant for the protein and complexes. In this system, the total concentrations of the activator, repressor, target and decoy sites are conserved:
@@
$T_{{\textsc{t}}}$ $=T+T_{A}+T_{R},$ $D_{{\textsc{t}}}=D+D_{A}+D_{R}.$ (8)
@@ III Paradoxical effects at steady state and transients
For the proposed synthetic circuit governed by equations (2)-(8), the paradoxical effect is captured in two scenarios. First, by varying the levels of activator in the circuit (by changing $u_{A}$) and second by varying the levels of decoy sites in the system. The knockout of the activator binding to the target is achieved by maintaining a high dissociation constant $k_{ta}=\tfrac{t_{a}^{-}}{t_{a}^{+}}$.
III-A Increasing activator causing unintended repression
@@
The proposed synthetic circuit exhibits the paradoxical effect given by:
[…]
where $\bar{Y}$ is the steady-state levels of the protein when other inputs ($u_{R}$) and parameters ($D_{{\textsc{t}}}$, $T_{{\textsc{t}}}$, reaction rate constants) are kept constant.
@@
$\overline{Y}=\frac{\beta\overline{T}+\kappa\overline{T_{A}}}{\gamma}$
where $\overline{x}$ is the steady-state concentration of $x$. At steady state, we have:
@@
where $K_{xy}=\tfrac{x_{y}^{-}+\bar{\delta}}{x_{y}^{+}}$. Note that as $k_{ta}\rightarrow\infty K_{ta}\rightarrow\infty$ for a finite $\bar{\delta}$. Substituting in equations (7)-(8), we get:
@@
$\therefore\overline{T}$ $=\frac{\overline{T_{{\textsc{t}}}}}{1+\frac{\overline{A}}{K_{ta}}+\frac{\overline{R}}{K_{tr}}},$ $\overline{D}=$ $\frac{\overline{D_{{\textsc{t}}}}}{1+\frac{\overline{A}}{K_{da}}+\frac{\overline{R}}{K_{dr}}}.$ (13)
@@
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=\frac{\beta}{\gamma}\frac{\mathrm{d}\overline{T}}{\mathrm{d}u_{A}}+\frac{\kappa}{\gamma}\frac{\mathrm{d}\overline{T}_{A}}{\mathrm{d}u_{A}}=\frac{\beta}{\gamma\bar{\delta}}\frac{\mathrm{d}\overline{T}}{\mathrm{d}A_{{\textsc{t}}}}+\frac{\kappa}{\gamma\bar{\delta}}\frac{\mathrm{d}\overline{T}_{A}}{\mathrm{d}A_{{\textsc{t}}}}.$
Applying product rule after substituting equation (9) in (13):
@@
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=\frac{1}{\gamma\bar{\delta}}$ $\left(\frac{\kappa\overline{T}}{K_{ta}}-\frac{\kappa\overline{A}+\beta K_{ta}}{K_{ta}}\frac{\overline{T}}{K_{ta}+\overline{A}+\frac{\overline{R}K_{ta}}{K_{tr}}}\right)\frac{\mathrm{d\overline{A}}}{\mathrm{d}A_{{\textsc{t}}}}$
$-\frac{\kappa\overline{A}+\beta K_{ta}}{K_{ta}}\frac{\overline{T}}{K_{tr}+\overline{R}+\frac{\overline{A}K_{tr}}{\gamma\bar{\delta}K_{ta}}}\frac{\mathrm{d\overline{R}}}{\mathrm{d}A_{{\textsc{t}}}}$ (14)
Calculating each derivative individually using equation (11) and substituting in equation (14) and then applying the limit of $K_{ta}\rightarrow\infty$, we get:
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=g_{1}(\overline{A},\overline{R})g_{2}(\overline{A},\overline{R})<0$ (15)
with
$g_{2}(\overline{A},\overline{R}):=\left[\frac{g_{3}(\overline{A},\overline{R})}{K_{ta}}-\frac{\beta\overline{R}D_{{\textsc{t}}}}{K_{tr}K_{da}K_{dr}[1+\frac{\overline{R}}{K_{dr}}+\frac{\overline{A}}{K_{da}}]^{2}}\right]$
[…]
Figure 7: Paradoxical effect in steady-state levels of the output protein portrayed for varying the levels of (a) activator input levels ($u_{A}$ for $D_{\textsc{t}}=20$) and (b) total decoy sites ($D_{{\textsc{t}}}$ for $u_{A}=10$) in the system. Two parameter bifurcation plots between $u_{A}$ and $D_{\textsc{t}}$ showing the existence of the paradoxical effects for (c) $K_{ta}\rightarrow\infty$ and its absence for (d) $K_{ta}=0.1$. The parameter values used for the simulation are: $K_{tr}=0.1,K_{dr}=0.1.K_{da}=0.1,T_{\textsc{t}}=1,R_{\textsc{t}}=10,\kappa=1000,\beta=100,\text{ and }\gamma=1$.
Figure 7 (a) shows the exhibition of paradoxical effect with increasing the input inducer levels of the activator. As $u_{A}$ is increased, the total amount of activators in the system increases. For low values of $K_{ta}$ ($K_{ta}=0.1$ and 1), the output protein level increases as expected with an increase in the levels of activator. Increasing $K_{ta}$ (to 10 and 100) and thereby gradually knocking out the direct influence of the activator on the target engenders the paradoxical effect, where increasing $u_{A}$ leads to an initial decrease in the output levels of $\overline{Y}$ followed by an increase. For the extreme case of complete knockout of the activator with $K_{ta}\rightarrow\infty$, we observe a monotonic decrease in the concentration of the protein. In the two-parameter bifurcation plots in Figure 7 (c), we see that for high $K_{ta}$ value, the paradoxical affect is observed only after a threshold value of $D_{\textsc{t}}$. For low values of $D_{\textsc{t}}$ (say $10^{0}$), increasing the activator levels have no significant effects on the output protein levels. On the other hand for high $D_{\textsc{t}}$ (say $10^{2}$), increasing the activator levels shows a decrease in the output protein levels hence the paradoxical effect. In the case of low $K_{ta}$ value in Figure 7 (d), we see that increasing $u_{A}$, increases the protein levels irrespective of the value of $D_{{\textsc{t}}}$.
@@
Proof.  The presence of basal expression is captured by $\beta$, transcription rate in the inactive/neutral form of the target. Substituting $\beta=0$ in equation (15):
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=g_{1}(A,R)\left[\frac{g_{3}(A,R)}{K_{ta}}\right]>0$
where $g_{3}(A,R)>0$ and $g_{1}(A,R)>0$ for all $A>0$ and $R>0$. Therefore, increasing the amount of activator causes an increase in the output levels irrespective of the value of $K_{ta}$.

### v2

@@ I INTRODUCTION AND BACKGROUND
The field of synthetic biology has as its ultimate goal to program new or modify existing biological systems, for applications ranging from cell therapies and regenerative medicine to biosensing and biofuel production. In general, a significant obstacle to the development of synthetic biological circuits is the influence of compositional context: the fundamental characteristics of a circuit alter in the presence of additional components due to competition for resources, off-target interactions, genetic context, growth rate feedback loops, and retroactivity effects. See e.g. [1, 2, 3, 4, 5, 6] and [7] for an overview. Unless one designs mathematically-validated control circuits to compensate for uncertainty, designers will need to re-adjust each part whenever new elements are integrated into a system.
@@
Since it is hard to test this mechanistic hypothesis through epigenetic modifications in mammalian cells, we propose here a synthetic biology analog involving transcription factors in bacterial cells. In this paper, we review the mechanism and make mathematical predictions from a model, as well as a proposed implementation using CRISPRi/a technology. Experimental work is ongoing. In the remainder of this section, we review our previous results [8].
@@ I-A Review of previous work
Motivation: Paradoxical gene activity pattern. In the aforementioned experiments, a paradoxical pattern of gene activation was observed that cannot be easily explained by known local gene regulatory interactions. Consider a gene of interest (e.g. ZEB1 which is a major driver of EMT). As shown in Figure 2,
@@
the gene is nominally repressed due to the dominant activity of the repressor (PRC2). However, when the repressor is knocked out, the gene achieves a mediocre amount of activation. In the second knockout experiment, an activating regulator (KMT2D) is knocked out. In that case, the gene achieves maximal activation despite the fact that its main repressor is not knocked-out. The main question is: how would the direct knockout of a repressing regulator be less effective in activating a gene compared to knocking out an activator? We summarize our answer next.
@@
    There is a large number of targets per regulator, e.g. hundreds or thousands.
@@
This simplified model can partially explain the paradoxical results shown in Figure 2. It shows how repressor knockout can fail to activate a target gene, and how an activator knockout can activate the target. However, it does not show how can an activator knockout be more effective at activation than a repressor knockout. This is since the mechanism of activation depends on the sequestration of the repressor, hence it cannot yield an activation that is stronger than a full repressor knockout. The paper [8] shows how to obtain the full spectrum of behaviors using three factors: one repressor and two activators of different strengths. On the other hand, one can obtain all three cases in Figure 2 provided that partial knock-outs (i.e., “knock-downs”) of genes are possible.
As explained in the introduction, it is very difficult to test this mechanistic explanation of the paradoxical effects through epigenetic modifications in mammalian cells. Thus, in this work, we describe and analyze a synthetic biology knock-down model with two factors that uses transcription factors in bacterial cells.
@@ II Proposed synthetic construct
The proposed circuit (shown in Figure 5) consists of a pool of shared resources (activators and repressors) regulating the production of the output protein by the target gene while competing with the rest of the genome, modeled as decoy sites.
@@
The transcription factors, activator ($A$, produced at a constant rate $u_{A}$) and repressor ($R$, produced at a constant rate $u_{R}$) bind to the target sites ($T$) to transform into an active form ($T_{A}$) and silenced form ($T_{R}$). The active form of the target undergoes transcription and translation to produce the output protein ($Y$) at a rate $\kappa>\beta$, where $\beta$ is the basal transcription rate in the inactive form of the target. The decoy sites ($D$) sequester the resources by forming $D_{A}$ and $D_{R}$ complexes respectively. The chemical reactions involved are:
@@
Here, $\gamma$ and $\bar{\delta}$ are the corresponding decay rates constant for the protein and complexes. We assume that transcription factors degrade and dilute irrespective of being bound to a gene. The corresponding reaction rate equations (RREs) can be obtained using mass action kinetics as [11]:
$\dot{A}$ $=u_{A}-d_{a}^{+}DA+d_{a}^{-}D_{A}-t_{a}^{+}TA+t_{a}^{-}T_{A}-\bar{\delta}A$ (2)
$\dot{R}$ $=u_{R}-d_{r}^{+}DR+d_{r}^{-}D_{R}-t_{r}^{+}TR+t_{r}^{-}T_{R}-\bar{\delta}R$ (3)
@@
We define the total concentrations of the activator, repressor, target, and decoy species as:
@@
$T_{{\textsc{t}}}$ $=T+T_{A}+T_{R},$ $D_{{\textsc{t}}}=D+D_{A}+D_{R}$ (10)
with their time derivatives being:
$\dot{A_{{\textsc{t}}}}$ $=u_{A}-\bar{\delta}A,$ $\dot{R_{{\textsc{t}}}}=u_{R}-\bar{\delta}R$ (11)
$\dot{T_{{\textsc{t}}}}$ $=0,$ $\dot{D_{{\textsc{t}}}}=0.$ (12)
@@ III Paradoxical effects at steady state and transients
For the proposed synthetic circuit governed by equations (2)-(10), the paradoxical effect is captured in two scenarios. First, by varying the levels of activator in the circuit (by changing $u_{A}$) and second by varying the levels of decoy sites in the system. The knockout of the activator binding to the target is achieved by maintaining a high dissociation constant $k_{ta}=\tfrac{t_{a}^{-}}{t_{a}^{+}}$. Note: hereon the steady-state concentration of the species $x$ is denoted as $\overline{x}$.
III-A Increasing activator causes unintended repression
@@
The output $\overline{Y}$ of our synthetic circuit depends in a complicated manner on the input $u_{A}$:
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=(\kappa-\beta)\frac{C_{1}(u_{A})}{k_{ta}}+C_{2}(u_{A}),$ (13)
where
$C_{1}$ $=C\left(1+\frac{\kappa\overline{R}}{(\kappa-\beta)k_{tr}}+\frac{\tfrac{T_{{\textsc{t}}}}{k_{tr}}}{1+\tfrac{\overline{R}}{k_{tr}}+\tfrac{\overline{A}}{k_{ta}}}+\frac{\tfrac{D_{{\textsc{t}}}}{k_{dr}}(1+\tfrac{\overline{A}}{k_{ta}})}{(1+\tfrac{\overline{R}}{k_{dr}}+\tfrac{\overline{A}}{k_{da}})^{2}}\right)$
$C_{2}$ $=\left[\frac{\kappa}{k_{ta}}-\frac{\beta}{k_{da}}\right]\frac{\overline{R}D_{{\textsc{t}}}C}{k_{tr}k_{dr}(1+\tfrac{\overline{R}}{k_{dr}}+\tfrac{\overline{A}}{k_{da}})^{2}}$
with $\overline{A}$ and $\overline{R}$ depend on $u_{A}$. The full expression for $C$ is provided in Appendix -A.
Remark 1
Notice that $C_{1}>0$ and
$C_{2}\begin{cases}>0&\text{if }k_{ta}<\frac{\kappa k_{da}}{\beta}\\ <0&\text{if }k_{ta}>\frac{\kappa k_{da}}{\beta}.\end{cases}$
This means that $\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}$ is the sum of a positive term and a term which is negative for appropriate parameter choices. In particular, for large values of $k_{ta}$ the negative second term dominates, so that the dependence of $\overline{Y}$ on $u_{A}$ is negative:
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}<0\quad\text{for}\quad k_{ta}\rightarrow\infty,$
contradicting the expectation that $A$ is an activator and hence larger values of $u_{A}$ should result in larger values of $\overline{Y}$.
@@
$\overline{Y}=\frac{\beta\overline{T}+\kappa\overline{T_{A}}}{\gamma}.$ (14)
At steady state, we have:
@@
where $K_{xy}=\tfrac{x_{y}^{-}+\bar{\delta}}{x_{y}^{+}}$. Note that as $k_{ta}\rightarrow\infty\implies K_{ta}\rightarrow\infty$ for a finite $\bar{\delta}$. Substituting in equations (9)-(10), we get:
@@
$\therefore\overline{T}$ $=\frac{\overline{T_{{\textsc{t}}}}}{O_{t}},$ $\overline{D}=$ $\frac{\overline{D_{{\textsc{t}}}}}{O_{d}}.$ (19)
where
$O_{t}$ $=1+\frac{\overline{A}}{K_{ta}}+\frac{\overline{R}}{K_{tr}},$ $O_{d}=$ $1+\frac{\overline{A}}{K_{da}}+\frac{\overline{R}}{K_{dr}}$
@@
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=\frac{\beta}{\gamma}\frac{\mathrm{d}\overline{T}}{\mathrm{d}u_{A}}+\frac{\kappa}{\gamma}\frac{\mathrm{d}\overline{T}_{A}}{\mathrm{d}u_{A}}=\frac{\beta}{\gamma\bar{\delta}}\frac{\mathrm{d}\overline{T}}{\mathrm{d}A_{{\textsc{t}}}}+\frac{\kappa}{\gamma\bar{\delta}}\frac{\mathrm{d}\overline{T}_{A}}{\mathrm{d}A_{{\textsc{t}}}},$
as $A_{{\textsc{t}}}=\tfrac{u_{A}}{\bar{\delta}}$ from equation (11). Applying product rule after substituting equation (15) in (19):
@@
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=\frac{1}{\gamma\bar{\delta}}$ $\left(\frac{\kappa\overline{T}}{K_{ta}}-\frac{\kappa\overline{A}+\beta K_{ta}}{K_{ta}}\frac{\overline{T}}{K_{ta}+\overline{A}+\frac{\overline{R}K_{ta}}{K_{tr}}}\right)\frac{\mathrm{d\overline{A}}}{\mathrm{d}A_{{\textsc{t}}}}-\frac{\kappa\overline{A}+\beta K_{ta}}{K_{ta}}\frac{\overline{T}}{K_{tr}+\overline{R}+\frac{\overline{A}K_{tr}}{\gamma\bar{\delta}K_{ta}}}\frac{\mathrm{d\overline{R}}}{\mathrm{d}A_{{\textsc{t}}}}$ (20)
Calculating each derivative individually using equation (17) and substituting in equation (20), we get:
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}u_{A}}=(\kappa-\beta)\frac{C_{1}}{k_{ta}}+C_{2}.$
Details of these computations are provided in Appendix -A.
Figure 6: Paradoxical effect in steady-state levels of the output protein portrayed for varying the levels of (a) activator input levels ($u_{A}$ for $D_{\textsc{t}}=20$) and (b) total decoy sites ($D_{{\textsc{t}}}$ for $u_{A}=10$) in the system. Two dimensional response plot between $u_{A}$ and $D_{\textsc{t}}$ showing the existence of the paradoxical effects for (c) $K_{ta}\rightarrow\infty$ and its absence for (d) $K_{ta}=0.1$. The parameter values used for the simulation are: $K_{tr}=0.1,K_{dr}=0.1,K_{da}=0.1,T_{\textsc{t}}=1,R_{\textsc{t}}=10,\kappa=1000,\beta=100,\text{ and }\gamma=1$.
Figure 6-a shows the exhibition of paradoxical effect with increasing the input inducer levels of the activator. As $u_{A}$ is increased, the total amount of activators in the system increases. For low values of $K_{ta}$ ($K_{ta}=0.1$ and 1), the output protein level increases as expected with an increase in the levels of activator. Increasing $K_{ta}$ (to 10 and 100) and thereby gradually knocking out the direct influence of the activator on the target engenders the paradoxical effect, where increasing $u_{A}$ leads to an initial decrease in the output levels of $\overline{Y}$ followed by an increase. For the extreme case of complete knockout of the activator with $K_{ta}\rightarrow\infty$, we observe a monotonic decrease in the concentration of the protein. In the two-parameter bifurcation plots in Figure 6-c, we see that for high $K_{ta}$ value, the paradoxical affect is observed only after a threshold value of $D_{\textsc{t}}$. For low values of $D_{\textsc{t}}$ (say $10^{0}$), increasing the activator levels have no significant effects on the output protein levels. On the other hand for high $D_{\textsc{t}}$ (say $10^{2}$), increasing the activator levels shows a decrease in the output protein levels hence the paradoxical effect. In the case of low $K_{ta}$ value in Figure 6-d, we see that increasing $u_{A}$, increases the protein levels irrespective of the value of $D_{{\textsc{t}}}$.
@@
Proof.  The presence of basal expression is captured by $\beta$, transcription rate in the inactive/neutral form of the target. Substituting $\beta=0$ in equation (13) implies $C_{2}>0\forall\ k_{ta}$. Therefore, increasing the amount of activator increases the output levels irrespective of the value of $k_{ta}$.